OpenStudy (anonymous):

Evaluate the expression if a = 2. 5a4 A. 30 B. 40 C. 60 D. 80

OpenStudy (charlotte123):

5a4? what is that for? :)

OpenStudy (charlotte123):

hello? D:

OpenStudy (texaschic101):

5a^4 when a = 2 5 * 2^4 = 5 * 16 = 80

OpenStudy (charlotte123):

so that's what it was xD 5a^4

OpenStudy (anonymous):

D. \[5a ^{4}\] 5(2)^4 2*2*2*2 5(16)=80.

OpenStudy (texaschic101):

yes..thats what it was